### v4

@@ 5.2 Research Returns to Scale in Cohort Size
Estimates of of a linear version of equation 9 using the full sample, reported in columns 1-2 and 5-6 of Panel A in Table LABEL:t:_CRTS, indicate that DI publications increase by $1.3-1.5$ per student, while T6 publications increase by $0.3-0.4$. These estimates are close to the mean post-publication statistics in Appendix Table LABEL:t:appendixstats. Estimates are similar when computed with and without school effects. Results using the sample of economics department graduates only, reported in Panel B, are also similar to those for the full sample. [^21] Cohort size in this case is the number of identifiable economics department graduates.
@@
The estimated marginal effect of cohort size in equation (9) is $\hat{\kappa}_{1}+2\hat{\kappa}_{2}n_{sc}$. The estimated change in returns to scale as a function of cohort size is therefore $2\hat{\kappa}_{2}$. Estimates of $\kappa_{2}$ for DI publications in the full sample imply a modest effect-change gradient of .0066(2×0.0033)-.0066(2$\times$-0.0033)- .0066 ( 2 × - 0.0033 ) when estimated without school effects. This flips to equally-modest-but positive $0.0042$ when school effects are added. Inclusion of school effects flips the sign of the other small estimated gradients elsewhere in the table as well: the implied change in returns to scale as enrollment grows are therefore neither large nor consistently negative.
@@ 6 Summary and Conclusions: Selection, Training, and Luck on the Path to PhD
The possibility of advisor-advisee sorting and the attendant selection bias motivates 2SLS analysis using dummies for cohort-by-school to instrument advising features. 2SLS estimation uncovers only weak evidence of causal effects of advising features on PhD student research success. In particular, at the level of cohorts and schools, only average advisor research predicts research success for the average graduate. 2SLS estimates of effects of past student success and research affinity are uniformly positive but mostly not significantly different from zero.
@@
Notes: This figure reports estimated cohort effects from a student-level Poisson regression of Deep Impact and Top 6 publication counts for years c+1 to c+6 on cohort and school effects. The reference year is 1989 and the sample starts with the 1989 cohort. Results for Deep Impact are on the left; results for Top 6 are on the right. Estimates in Panel A use the sample of economics and related program graduates (dependent variable means are 1.53 for Deep Impact publication counts and 0.34 for Top 6 publication counts). Panel B reports estimates for identifiable economics department graduates only (dependent variable means are 1.59 for Deep Impact publication counts and 0.36 for Top 6 publications). Standard errors are clustered on school-by-year.
@@
Figure 13: Student Success and No. of Advisees for 235 Advising-Loaded Advisors
@@
Notes: The left panel shows the number of graduates advised by advisors who change affiliations, with separate counts for cohorts graduating at the former and new affiliation before and after a transition year. The right panel plots the number of graduates citing a transitioning advisors’ work at the previous and new affiliation. The analysis includes 367 fixed-super advisors (ranked by advisee count) with at least one affiliation change, as indicated by the affiliation variable on their EconLit publications. The sample is restricted to research-active advisors whose affiliations are observed in at least one-third of their research-active years (the years between their first and last publications). Advisors in this sample advised at least one student as of an affiliation change.
@@
Notes: This figure plots overall school-by-cohort averages of Deep Impact and Top 6 publication output in years $c+1$ to $c+6$ for graduates in cohort $c$. As in the previous figure, these are computed separately for schools losing or gaining a transitioning advisor, looking at cohorts completing degrees before and after a transition year. Transitioners here are 385 fixed-duper advisors (ranked by past student success) with at least one affiliation change, as indicated by the affiliation variable on their EconLit publications, limited to research-active advisors whose affiliations are observed in more than one-third of their research-active years, with at least one advisee graduating ahead of the affiliation change. Publication statistics are for all graduates from affected schools.
@@
Berkeley 854 35.6 .45 103 4.3 .34
Chicago 681 28.4 .37 105 4.4 .52
Harvard 792 33.0 .53 91 3.8 .41
MIT 633 26.4 .56 151 6.3 .30
Northwestern 485 20.2 .53 54 2.3 .48
Princeton 467 19.50 .56 27 1.1 .44
Stanford 532 22.2 .47 215 9 .44
Yale 454 18.9 .50 52 2.2 .29
All 4898 25.5 .49 798 4.2 .40
@@ A. Research Productivity
Super Advisor Research 0.69∗∗∗ 0.21∗∗∗ 0.48∗∗∗ 0.68∗∗∗
@@ B. Advising Load
Super Advising Load 0.43∗∗∗ 0.17∗∗∗ 0.32∗∗∗ 0.69∗∗∗
@@ C. Successful Students
Past Student Success 0.43∗∗∗ 0.14∗∗∗ 0.26∗∗∗ 0.36∗∗∗
@@
Super Past Student Success 0.56∗∗∗ 0.18∗∗∗ 0.36∗∗∗ 0.51∗∗∗
@@
Dep. var. mean 1.51 0.34 1.51 0.34
Notes: The dependent variable in columns 1, 2, 5, and 6 is student Deep Impact publications in the six years after graduation ($c+1$ to $c+6$ for cohort $c$). The dependent variable in columns 3, 4, 7, and 8 sums top-6 publications. Research productivity regressors average a cohort-$c$ graduate’s advisors’ Deep Impact publications in the five years before $c$. Advising load regressors average a graduate’s advisors’ advisee counts in the five years before $c$. For a given graduate, the successful student regressor averages the number of Deep Impact publications by a graduate’s advisors’ past advisees as described by equation 4 in the text. “Super” regressors are dummies indicating whether at least one of a student’s advisors was among the 10% most prolific advisors relative to all other advisors who advised students graduating in the five years preceding the student’s graduation year. Models for each panel are run separately. All models control for school and cohort effects, graduates’ advisor team size, graduate gender, unclassified gender, missing thesis PDF, and an economics department dummy. Estimates are for the economics + related sample of 5682 students that graduated between 1994 and 2017. Standard errors are clustered on school-by-cohort (192 clusters for 8 schools $\times$ 24 cohorts).
@@
Notes: Dependent variables and sample are as described in the note to Table 2. Models control for school and cohort effects, graduates’ advisor team size, graduate gender, unclassified gender, missing thesis PDF, and an economics department dummy. Standard errors are clustered on school-by-cohort (192 clusters). Of the 5682 economics+related program graduates in the 1994-2017 cohorts, 547 published by the end of graduation year. Of those, 187 coauthored with an advisor and 58 with a classmate. Cites Advisor in Thesis is a dummy variable indicating PhD theses that cite an advisor’s work.
Notes: This table ranks prolific advisors (who advised at least 30 graduates) by the research success of their graduate student advisees who earned degrees at one of the eight schools in our sample in 1989-2023, truncated to advisors whose past student success is greater than one. Student success averages post-graduation publications (i.e., in years $c+1$ to $c+6$ for a graduate in cohort $c$) over an advisor’s students. The affiliation column in the table indicates the modal school from an advisor’s advisees graduated. The next-to-last column in the table ranks advisors by number of advisees. Rankings change little when student publications exclude work coauthored with advisors.

### v5

@@ 5.2 Research Returns to Scale in Cohort Size
Estimates of of a linear version of equation 9 using the full sample, reported in columns 1-2 and 5-6 of Panel A in Table LABEL:t:_CRTS, indicate that DI publications increase by $1.4-1.6$ per student, while T6 publications increase by around $0.4$. These estimates are close to the mean post-publication statistics in Appendix Table LABEL:t:appendixstats. Estimates are similar when computed with and without school effects. Results using the sample of economics department graduates only, reported in Panel B, are also similar to those for the full sample. [^21] Cohort size in this case is the number of identifiable economics department graduates.
@@
The estimated marginal effect of cohort size in equation (9) is $\hat{\kappa}_{1}+2\hat{\kappa}_{2}n_{sc}$. The estimated change in returns to scale as a function of cohort size is therefore $2\hat{\kappa}_{2}$. Estimates of $\kappa_{2}$ for DI publications in the full sample imply a modest effect-change gradient of .0132(2×0.0066)-.0132(2$\times$-0.0066)- .0132 ( 2 × - 0.0066 ) when estimated without school effects. This reduces to .$-.003$ when school effects are added. The implied change in returns to scale as enrollment grows are therefore very small.
@@ 6 Summary and Conclusions: Selection, Training, and Luck on the Path to PhD
The possibility of advisor-advisee sorting and the attendant selection bias motivates 2SLS analysis using dummies for cohort-by-school to instrument advising features. 2SLS estimation uncovers only weak evidence of causal effects of advising features on PhD student research success. In particular, at the level of cohorts and schools, only average advisor research predicts research success for the average graduate. 2SLS estimates of effects of past student success and research affinity are uniformly positive but not significantly different from zero.
@@
Notes: This figure reports estimated cohort effects from a student-level Poisson regression of Deep Impact and Top 6 publication counts for years c+1 to c+6 on cohort and school effects. The reference year is 1989 and the sample starts with the 1989 cohort. Results for Deep Impact are on the left; results for Top 6 are on the right. Estimates in Panel A use the sample of economics and related program graduates (dependent variable means are 1.59 for Deep Impact publication counts and 0.39 for Top 6 publication counts). Panel B reports estimates for identifiable economics department graduates only (dependent variable means are 1.65 for Deep Impact publication counts and 0.42 for Top 6 publications). Standard errors are clustered on school-by-year.
@@
Figure 13: Student Success and No. of Advisees for 236 Advising-Loaded Advisors
@@
Notes: The left panel shows the number of graduates advised by advisors who change affiliations, with separate counts for cohorts graduating at the former and new affiliation before and after a transition year. The right panel plots the number of graduates citing a transitioning advisors’ work at the previous and new affiliation. The analysis includes 364 fixed-super advisors (ranked by advisee count) with at least one affiliation change, as indicated by the affiliation variable on their EconLit publications. The sample is restricted to research-active advisors whose affiliations are observed in at least one-third of their research-active years (the years between their first and last publications). Advisors in this sample advised at least one student as of an affiliation change.
@@
Notes: This figure plots overall school-by-cohort averages of Deep Impact and Top 6 publication output in years $c+1$ to $c+6$ for graduates in cohort $c$. As in the previous figure, these are computed separately for schools losing or gaining a transitioning advisor, looking at cohorts completing degrees before and after a transition year. Transitioners here are 388 fixed-duper advisors (ranked by past student success) with at least one affiliation change, as indicated by the affiliation variable on their EconLit publications, limited to research-active advisors whose affiliations are observed in more than one-third of their research-active years, with at least one advisee graduating ahead of the affiliation change. Publication statistics are for all graduates from affected schools.
@@
Berkeley 854 35.6 .46 103 4.3 .36
Chicago 681 28.4 .38 105 4.4 .53
Harvard 791 33.0 .54 91 3.8 .42
MIT 631 26.3 .58 150 6.3 .31
Northwestern 485 20.2 .54 54 2.3 .50
Princeton 467 19.5 .57 27 1.1 .44
Stanford 530 22.1 .48 216 9.0 .46
Yale 454 18.9 .52 52 2.2 .29
All 4893 25.5 .50 798 4.2 .41
@@ A. Research Productivity
Super Advisor Research 0.70∗∗∗ 0.24∗∗∗ 0.47∗∗∗ 0.67∗∗∗
@@ B. Advising Load
Super Advising Load 0.47∗∗∗ 0.19∗∗∗ 0.34∗∗∗ 0.70∗∗∗
@@ C. Successful Students
Past Student Success 0.43∗∗∗ 0.15∗∗∗ 0.25∗∗∗ 0.35∗∗∗
@@
Super Past Student Success 0.57∗∗∗ 0.20∗∗∗ 0.36∗∗∗ 0.49∗∗∗
@@
Dep. var. mean 1.56 0.39 1.56 0.39
Notes: The dependent variable in columns 1, 2, 5, and 6 is student Deep Impact publications in the six years after graduation ($c+1$ to $c+6$ for cohort $c$). The dependent variable in columns 3, 4, 7, and 8 sums top-6 publications. Research productivity regressors average a cohort-$c$ graduate’s advisors’ Deep Impact publications in the five years before $c$. Advising load regressors average a graduate’s advisors’ advisee counts in the five years before $c$. For a given graduate, the successful student regressor averages the number of Deep Impact publications by a graduate’s advisors’ past advisees as described by equation 4 in the text. “Super” regressors are dummies indicating whether at least one of a student’s advisors was among the 10% most prolific advisors relative to all other advisors who advised students graduating in the five years preceding the student’s graduation year. Models for each panel are run separately. All models control for school and cohort effects, graduates’ advisor team size, graduate gender, unclassified gender, missing thesis PDF, and an economics department dummy. Estimates are for the economics + related sample of 5676 students that graduated between 1994 and 2017. Standard errors are clustered on school-by-cohort (192 clusters for 8 schools $\times$ 24 cohorts).
@@
Notes: Dependent variables and sample are as described in the note to Table 2. Models control for school and cohort effects, graduates’ advisor team size, graduate gender, unclassified gender, missing thesis PDF, and an economics department dummy. Standard errors are clustered on school-by-cohort (192 clusters). Of the 5676 economics+related program graduates in the 1994-2017 cohorts, 575 published by the end of graduation year. Of those, 193 coauthored with an advisor and 61 with a classmate. Cites Advisor in Thesis is a dummy variable indicating PhD theses that cite an advisor’s work.
Notes: This table ranks prolific advisors (who advised at least 30 graduates) by the average DI publications of their advisees who earned degrees at one of the eight schools in our sample in 1989-2023, truncated to advisors for whom their students’ average DI publications exceed one. Row order ranks advisors by their students’ average post-graduation DI publications (i.e., in years $c+1$ to $c+6$ for a graduate in cohort $c$). The table also reports ranks by students’ average T6 publications, the number of students advised, and advisor rank by number advised. The affiliation column in the table indicates the modal school from which an advisor’s advisees graduated. Rankings change little when student publications exclude work coauthored with advisors.